OpenStudy (anonymous):

Amistre64, it's me again. Need help with an exponent again!

OpenStudy (amistre64):

didja try yelling at it?

OpenStudy (anonymous):

jI even grounded it and took away it's cell phone and it won't cooperate! I can't understand...it works with my 15 year old daughter!

OpenStudy (anonymous):

\[-3/2\times 2^{-2/9}\]

OpenStudy (anonymous):

dont punish them it makes it worse. They now think math is bad if you truly want to help arrange extra help and talk to the teacher

OpenStudy (anonymous):

Did you see my expression I need to solve?

OpenStudy (amistre64):

i see it

OpenStudy (amistre64):

we wanna derive it?

OpenStudy (anonymous):

No I need to figure out what it's = to.

OpenStudy (anonymous):

2^-2/9 equals what?

OpenStudy (amistre64):

\[\frac{-3}{2*2^{2/9}}\]

OpenStudy (amistre64):

\[\frac{-3}{2^{(9/9 + 2/9)}}\]

OpenStudy (amistre64):

we good so far?

OpenStudy (anonymous):

yes, I got that far. I know that answer, but what i got wasn't right according to them Like they know any better! Who are they to tell me I am always wrong!?

OpenStudy (anonymous):

; )

OpenStudy (amistre64):

:) \[\frac{-3}{2^{11/9}}\] \[\frac{-3}{\sqrt[9]{2^{11}}}\]

OpenStudy (amistre64):

\[2^{11} \iff 2^{9}*2^{2}\] \[\sqrt[9]{2^9.2^2}\iff \sqrt[9]{2^9}.\sqrt[9]{2^2}\] \[\frac{-3}{2\sqrt[9]{2^2}}\]

OpenStudy (amistre64):

\[\sqrt[9]{2^4}*\sqrt[9]{2^5} = \sqrt[9]{2^9}\iff 2\] right?

OpenStudy (anonymous):

yes, I remember that one from the s thing earlier! Yay me!

OpenStudy (amistre64):

\[\frac{-3}{2\sqrt[9]{2^4}}*\frac{\sqrt[9]{2^5}}{\sqrt[9]{2^5}} = \frac{-3\sqrt[9]{64}}{4}\]

OpenStudy (amistre64):

i typoed it dint i lol

OpenStudy (amistre64):

9rt(2^2) not equal 9rt(2^4) lol

OpenStudy (anonymous):

Actually we lucked out cuz they are leaving the radical in the denominator! The answer to the whole problem is cbrt 2 - 3/2ninth root of 4, just what you said! I am not surprised you got it right!

OpenStudy (amistre64):

yay!! i was dreading having to retype thaqt lol

OpenStudy (anonymous):

\[\sqrt[3]{2}- 3/2\sqrt[9]{4}\]

OpenStudy (anonymous):

Again and again and again, thank you.

OpenStudy (amistre64):

yw :)